# Transferable Reinforcement Learning via Generalized Occupancy Models

### v2

@@ 4.2 Transfer with Generalized Occupancy Models
where $\epsilon>0$ is a tuning parameter to ensure sufficient coverage for $\psi$. This suggests that the optimal outcome $\psi^{*}$ is the one that provides the highest future sum of rewards $w_{r}^{T}\psi^{*}$ that is valid under the environment dynamics and dataset coverage.
Figure 3: Planning with GOMs. Once a GOM is learned, the optimal action can be obtained by performing reward regression and searching for the optimal outcome under the dynamics to decode via the policy.
This optimization problem can be solved in a number of ways. Perhaps the most straightforward is via a random shooting technique [52], which samples a set of $\psi$ from $p(\psi|s)$ and chooses the one with the highest $w_{r}^{T}\psi.$ Sec. 5 bases our theoretical analysis on this planning method. Sec. 4.3 shows that for outcome models instantiated with specific model classes such as diffusion models, the planning problem can be simplified to guided sampling.
Once $\psi^{*}$ has been obtained, the action to execute in the environment can be acquired via the readout policy $\pi_{\rho}(a|s,\psi^{*})$. Fig. 3 shows the full the planning procedure, and we refer the reader to Appendix. F for the pseudocode. As described previously, GOMs enable rapid transfer to arbitrary new rewards in an environment without accumulating compounding error or requiring expensive test-time policy optimization. In this way, they can be considered a new class of models of transition dynamics that avoids the typical challenges in model-based RL and successor features.
@@ 4.3 Practical Instantiation
In this section, we provide a practical instantiation of GOMs that is used throughout our experimental evaluation. The first step to instantiate GOMs is to choose an expressive state feature that linearly expresses a broad class of rewards. We choose the state feature $\phi$ to be $d$-dimensional random Fourier features [51]. Next, the model class must account for the multimodal nature of outcomes and actions since multiple outcomes can follow from a state, and multiple actions can be taken to realize an outcome. To this end, we parametrize both the outcome model $p(\psi|s)$ and the readout policy $\pi(a|s,\psi)$ using a conditional diffusion model [22, 46]. We then train these models (optimize Equation 1,  2) by denoising score matching, a surrogate of maximum likelihood training [47].
Remarkably, when $p(\psi|s)$ is parameterized by a diffusion model, the special structure of the planning problem in Eq. 3 allows a simple variant of guided diffusion [13, 24] to be used for task-directed planning. In particular, taking the log of both sides of the constraint and recasting the constrained optimization via the penalty method, we get a penalized objective $\mathcal{L}=w_{r}^{T}\psi+\alpha(\log p(\psi|s)-\log\epsilon)$. Taking the gradient yields $\nabla_{\psi}\mathcal{L}(\psi,\alpha)=w_{r}+\alpha\nabla_{\psi}\log p(\psi|s).$ The expression for $\nabla_{\psi}\mathcal{L}(\psi,\alpha)$ is simply the score function $\nabla_{\psi}\log p(\psi|s)$ in standard diffusion training (Section 3), with the linear weights $w_{r}$ added as a guidance term. Planning then becomes doing stochastic gradient Langevin dynamics [56] to obtain an optimal $\psi^{*}$ sample, using $\nabla_{\psi}\mathcal{L}(\psi,\alpha)$ as the gradient. Guided diffusion removes the need for sampling a set of particles. As we show in Appendix E, it matches the performance of random shooting while taking significantly less inference time. In Appendix B we show that the guided diffusion procedure can alternatively be viewed as taking actions conditioned on a soft optimality variable.
5 Theoretical Analysis of GOMs
To provide a theoretical understanding of GOMs, we conduct an error analysis to connect the error in estimating the ground truth $p_{0}(\psi\mid s)$ to the suboptimality of the GOM policy, and then study when the GOM policy becomes optimal. We start our analysis conditioning on $\epsilon>0$ estimation error in the ground truth outcome distribution $p_{0}(\psi\mid s)$.
@@ 5 Theoretical Analysis of GOMs
Since GOMs capture the outcome distribution of the behavior policy $\pi_{\beta}$, we need a definition to evaluate a policy $\pi$ with respect to $\pi_{\beta}$.
@@
We proceed to use Definition 5.2 to characterize the suboptimality of the GOM policy. Let $\tau$ denote the sampling optimality of the random shooting planner in Sec. 4.2. Specifically, we expect to sample a top $\tau$ outcome $\psi$ from the behavior policy in $O(\frac{1}{\tau})$ samples, where $\mathbb{P}_{\pi_{\beta}}[w_{r}^{T}\psi\leq\sum_{t}\gamma^{t-1}r(s_{t})]\leq\tau$. The following result characterizes the suboptimality of the GOM policy. The proof is deferred to Appendix. A.
@@
Lastly, we extend our main theoretical result to the standard full data coverage condition in the offline RL literature, where the dataset contains all transitions [50, 59, 43]. The following theorem states that GOMs can output the optimal policy in this case. The proof is deferred to Appendix A.
@@
In deterministic MDPs, when $|\mathcal{A}\times\mathcal{S}|<\infty$, and $\forall(s,a)\in\mathcal{S}\times\mathcal{A},N(s,a,\mathcal{T}(s,a))\geq 1$, GOM is guaranteed to identify an optimal policy.
@@ 6 Experimental Evaluation
In our experimental evaluation, we aim to answer the following research questions. (1) Can GOMs transfer across tasks without expensive test-time policy optimization? (2) Can GOMs avoid the challenge of compounding error present in MBRL? (3) Can GOMs solve tasks with arbitrary rewards beyond goal-reaching problems? (4) Can GOMs go beyond the offline dataset, and accomplish “trajectory-stitching" to actualize outcomes that combine different trajectories?
@@
Figure 4: Evaluation domains: (1) D4RL Antmaze [14] (2) Franka Kitchen [14] (3) Preference-Based Antmaze with the goal of taking a particular path (4) Roboverse [45] robotic manipulation.
@@ 6.1 Problem Domains and Datasets
D4RL Antmaze D4RL Antmaze [14] is a navigation domain that involves controlling a quadruped to reach some designated goal location. Each task corresponds to reaching a different goal location. We use the D4RL dataset for pretraining and dense rewards described in Appendix D for adaptation.
Franka Kitchen Franka Kitchen [14] is a manipulation domain where the goal is to control a Franka arm to interact with appliances in the kitchen. Each task corresponds to interacting with a set of items. We use the D4RL dataset for pretraining and standard sparse rewards for adaptation.
Preference Antmaze Preference Antmaze is a variant of D4RL Antmaze [14] where the goal is to reach the top right corner starting from the bottom left corner. The two tasks in this environment correspond to the two paths to reach the goal, simulating human preferences. We collect a custom dataset and design reward functions for each preference.
Roboverse Robotic Manipulation Roboverse [45] is a tabletop manipulation environment with a robotic arm completing multi-step problems. Each task consists of two phases, and the offline dataset contains separate trajectories of each phase but not full task completion. A sparse reward is assigned to each time step of task completion.
Table 1: Offline multitask RL on AntMaze and Kitchen. GOMs show superior transfer performance (in average episodic return) than successor features, model-based RL, and misspecified goal-conditioned baselines.
GOM (Ours) USF FB RaMP MOPO COMBO GC-IQL
@@ 6.2 Baseline Comparisons
Successor Features We compare with three methods from the successor feature line of work. Universal SF (USF) [6] overcomes the policy dependence of SF by randomly sampling reward weights $z$ and jointly learning a successor feature predictor $\psi_{z}$ and a policy $\pi_{z}$ conditioned on $z$. $\psi_{z}$ captures the successor feature of $\pi_{z}$, while $\pi_{z}$ is trained to maximize the reward described by $z$. Forward-Backward Representation (FB) [53, 54] follows the same paradigm but jointly learns a feature network by parameterizing the successor measure as an inner product between a forward and a backward representation. RaMP [8] removes the policy dependence of SF by predicting cumulative features from an initial state and an open-loop sequence of actions, which can be used for planning.
@@
6.3 Do GOMs learn transferable behavior across tasks?
@@
Figure 5: Transfer across tasks with GOMs and COMBO [62] in medium antmaze. Each tile corresponds to a different task, with color of the tile indicating the normalized return. GOMs successfully transfer across a majority of tasks, while MBRL [62] struggles on tasks that are further away from the initial location.
@@
We evaluate GOMs on transfer problems, where the dynamics are shared, but the reward functions vary. We train GOMs using the data distributions provided with the standard D4RL datasets  [14]. While GOMs in principle can identify the task reward from a small number of state-reward pairs, we relabel the offline dataset with test-time rewards to remove the confounding factor of exploration.
@@
Table 1 reports the episodic return of transferring to the default D4RL tasks. GOMs are able to adapt to new tasks with a simple linear regression, showing significantly higher transfer performance than successor features (mismatch between training and evaluation policy sets), model-based RL (compounding error) and goal-conditioned RL (goal distribution misspecification). Notably, we show in Appendix E that GOMs are even competitive with goal-conditioned RL methods trained on test-time goals. The transferability of GOMs can also be seen in Fig  5, where we plot the performance of GOMs across various tasks (corresponding to different tiles in the maze). We see that GOMs have less degradation across tasks than model-based RL  [62].
@@
Although the GOM framework and theoretical results are derived under deterministic MDPs, we emphasize that the D4RL antmaze datasets are collected with action noise, emulating stochastic transitions. These results indicate that GOM are practically applicable to some range of stochastic settings, although we expect it to perform better in purely deterministic settings.
6.4 Can GOMs solve tasks with arbitrary rewards?
@@
While methods like goal-conditioned RL [39, 16] are restricted to shortest path goal-reaching problems, GOMs are able to solve problems with arbitrary reward functions. This is crucial when the reward is not easily reduced to a particular “goal". To validate this, we evaluate GOMs on tasks that encode nontrivial human preferences in a reward function, such as particular path preferences in antmaze. In this case, we have different rewards that guide the agent specifically down the path to the left and the right, as shown in Fig 4. As we see in Table 3, GOMs and model-based RL obtain policies that respect human preferences and are performant for various rewards. Goal-conditioned algorithms are unable to disambiguate preferences and end up with some probability of taking each path.
@@
Table 2: Evaluation on non-goal-conditioned tasks. GOMs are able to solve non-goal-conditioned tasks, taking different paths in preference antmaze (Fig 4), while goal-conditioned RL cannot optimize for arbitrary rewards.
GOM (Ours) COMBO GC-IQL
@@
Table 3: Evaluation of trajectory stitching ability of GOMs. GOMs outperform non-stitching baselines, demonstrating their abilities to recombine outcomes across trajectory segments
GOM (Ours) RaMP DT
@@
6.5 Do GOMs perform trajectory stitching?
The ability to recover optimal behavior by combining suboptimal trajectories, or “trajectory stitching," is crucial to off-policy RL methods as it ensures data efficiency and avoids requirements for exponential data coverage. GOMs naturally enables this type of trajectory stitching via the distributional Bellman backup, recovering “best-in-data” policies for downstream tasks. To evaluate the ability of GOMs to perform trajectory stitching, we consider the environments introduced in  [45]. Here, the data only consists of trajectories that complete individual subtasks (e.g. grasping or placing), while the task of interest rewards the completion of both subtasks. Since the goal of this experiment is to evaluate stitching, not transfer, we choose the features as the task rewards $\phi(s)=r(s)$. We find that GOMs are able to show non-trivial success rates by stitching together subtrajectories. Since RaMP [8] predicts the summed features from a sequence of actions, and the optimal action sequence is not present in the dataset, it fails to solve any task. Likewise, return-conditioned supervised learning methods like Decision Transformer [9] do not stitch together trajectories and fails to learn meaningful behaviors.

### v3

@@ 4.2 Zero-Shot Policy Optimization with Distributional Successor Features
where $\epsilon>0$ is a hyperparameter to ensure sufficient coverage for $\psi$. This suggests that the optimal outcome $\psi^{*}$ is the one that provides the highest future sum of rewards $w_{r}^{T}\psi^{*}$ while being valid under the environment dynamics and dataset coverage.
Figure 3: Zero-shot policy optimization with DiSPOs. Once a DiSPO is learned, the optimal action can be obtained by performing reward regression and searching for the optimal outcome under the dynamics to decode via the policy.
This optimization problem can be solved in a number of ways. The most straightforward is via random shooting [53], which samples a set of $\psi$ from $p(\psi|s)$ and chooses the one with the highest $w_{r}^{T}\psi.$ Sec. 5 bases our theoretical analysis on this technique. Sec. 4.3 shows that for outcome models instantiated with diffusion models, the optimization problem can be simplified to guided diffusion sampling.
Once $\psi^{*}$ has been obtained, the action to execute in the environment can be acquired via the readout policy $\pi_{\rho}(a|s,\psi^{*})$. Fig. 3 shows the full policy optimization procedure, and we refer the reader to Appendix. F for the pseudocode. As described previously, DiSPOs enable zero-shot transfer to arbitrary new rewards in an environment without accumulating compounding error or requiring expensive test-time policy optimization. In this way, they can be considered a new class of models of transition dynamics that avoids the typical challenges in model-based RL and successor features.
@@ 4.3 Practical Instantiation
In this section, we provide a practical instantiation of DiSPOs that is used throughout our experimental evaluation. The first step to instantiate DiSPOs is to choose an expressive state feature that linearly expresses a broad class of rewards. We choose the state feature $\phi$ to be $d$-dimensional random Fourier features [52]. Next, the model class must account for the multimodal nature of outcomes and actions since multiple outcomes can follow from a state, and multiple actions can be taken to realize an outcome. To this end, we parametrize both the outcome model $p(\psi|s)$ and the readout policy $\pi(a|s,\psi)$ using a conditional diffusion model [23, 47]. We then train these models (optimize Equation 1,  2) by denoising score matching, a surrogate of maximum likelihood training [48].
Remarkably, when $p(\psi|s)$ is parameterized by a diffusion model, the special structure of the optimization problem in Eq. 3 allows a simple variant of guided diffusion [14, 25] to be used for policy optimization. In particular, taking the log of both sides of the constraint and recasting the constrained optimization via the penalty method, we get a penalized objective $\mathcal{L}=w_{r}^{T}\psi+\alpha(\log p(\psi|s)-\log\epsilon)$. Taking the gradient yields
$\nabla_{\psi}\mathcal{L}(\psi,\alpha)=w_{r}+\alpha\nabla_{\psi}\log p(\psi|s).$ (4)
The expression for $\nabla_{\psi}\mathcal{L}(\psi,\alpha)$ is simply the score function $\nabla_{\psi}\log p(\psi|s)$ in standard diffusion training (Section 3), with the linear weights $w_{r}$ added as a guidance term. Planning then becomes doing stochastic gradient Langevin dynamics [57] to obtain an optimal $\psi^{*}$ sample, using $\nabla_{\psi}\mathcal{L}(\psi,\alpha)$ as the gradient. Guided diffusion removes the need for sampling a set of particles. As shown in Appendix E, it matches the performance of random shooting while taking significantly less inference time. In Appendix B, we show that the guided diffusion procedure can alternatively be viewed as taking actions conditioned on a soft optimality variable.
5 Theoretical Analysis of Distributional Successor Features for Zero-Shot Policy Optimization
To provide a theoretical understanding of DiSPOs, we conduct an error analysis to connect the error in estimating the ground truth $p_{0}(\psi\mid s)$ to the suboptimality of the DiSPO policy, and then study when the DiSPO policy becomes optimal. We start our analysis conditioning on $\epsilon>0$ estimation error in the ground truth outcome distribution $p_{0}(\psi\mid s)$.
@@ 5 Theoretical Analysis of Distributional Successor Features for Zero-Shot Policy Optimization
Since DiSPOs capture the outcome distribution of the behavior policy $\pi_{\beta}$, we need a definition to evaluate a policy $\pi$ with respect to $\pi_{\beta}$.
@@
We proceed to use Definition 5.2 to characterize the suboptimality of the DiSPO policy. Let $\tau$ denote the sampling optimality of the random shooting planner in Sec. 4.2. Specifically, we expect to sample a top $\tau$ outcome $\psi$ from the behavior policy in $O(\frac{1}{\tau})$ samples, where $\mathbb{P}_{\pi_{\beta}}[w_{r}^{T}\psi\leq\sum_{t}\gamma^{t-1}r(s_{t})]\leq\tau$. The following result characterizes the suboptimality of the DiSPO policy. The proof is deferred to Appendix. A.
@@
Lastly, we extend our main theoretical result to the standard full data coverage condition in the offline RL literature, where the dataset contains all transitions [51, 60, 44]. The following theorem states that DiSPOs can output the optimal policy in this case. The proof is deferred to Appendix A.
@@
In deterministic MDPs, when $|\mathcal{A}\times\mathcal{S}|<\infty$, and $\forall(s,a)\in\mathcal{S}\times\mathcal{A},N(s,a,\mathcal{T}(s,a))\geq 1$, DiSPOs are guaranteed to identify an optimal policy.
@@ 6 Experimental Evaluation
In our experimental evaluation, we aim to answer the following research questions. (1) Can DiSPOs transfer across tasks without expensive test-time policy optimization? (2) Can DiSPOs avoid the challenge of compounding error present in model-based RL? (3) Can DiSPOs solve tasks with arbitrary rewards beyond goal-reaching problems? (4) Can DiSPOs go beyond the offline dataset, and accomplish “trajectory-stitching" to actualize outcomes that combine different subtrajectories?
@@
Figure 4: Evaluation domains: (1) D4RL Antmaze [15] (2) Franka Kitchen [15] (3) Hopper  [9] (4) Preference-Based Antmaze with the goal of taking a particular path (5) Roboverse [46] robotic manipulation.
@@ 6.1 Problem Domains and Datasets
Antmaze [15] is a navigation domain that involves controlling a quadruped to reach some designated goal location. Each task corresponds to reaching a different goal location. We use the D4RL dataset for pretraining and dense rewards described in Appendix D for adaptation.
Franka Kitchen [15] is a manipulation domain where the goal is to control a Franka arm to interact with appliances in the kitchen. Each task corresponds to interacting with a set of items. We use the D4RL dataset for pretraining and standard sparse rewards for adaptation.
Hopper [7, 9] is a locomotion domain that involves controlling a hopper to perform various tasks, including hopping forward, hopping backward, standing, and jumping. We use the offline dataset from [9] for pretraining and shaped rewards for adaptation.
Preference Antmaze is a variant of D4RL Antmaze [15] where the goal is to reach the top right corner starting from the bottom left corner. The two tasks in this environment correspond to the two paths to reach the goal, simulating human preferences. We collect a custom dataset and design reward functions for each preference.
Roboverse [46] is a tabletop manipulation environment with a robotic arm completing multi-step problems. Each task consists of two phases, and the offline dataset contains separate trajectories of each phase but not full task completion. A sparse reward is assigned to each time step of task completion.
Table 1: Offline multitask RL on AntMaze and Kitchen. DiSPOs show superior transfer performance (in average episodic return) than successor features, model-based RL, and misspecified goal-conditioned baselines.
DiSPO (Ours) USFA FB RaMP MOPO COMBO GC-IQL
@@
hopper-forward 566 $\pm$ 63 487 $\pm$ 110 452 $\pm$ 59 470 $\pm$ 16 493 $\pm$ 114 982 $\pm$ 157 -
hopper-backward 367 $\pm$ 15 261 $\pm$ 68 269 $\pm$ 77 220 $\pm$ 15 596 $\pm$ 211 194 $\pm$ 74 -
hopper-stand 800 $\pm$ 0 685 $\pm$ 130 670 $\pm$ 120 255 $\pm$ 15 800 $\pm$ 0 600 $\pm$ 111 -
hopper-jump 832 $\pm$ 22 746 $\pm$ 112 726 $\pm$ 35 652 $\pm$ 28 753 $\pm$ 51 670 $\pm$ 109 -
@@ 6.2 Baseline Comparisons
Successor Features We compare with three methods from the successor feature line of work. USFA [6] overcomes the policy dependence of SF by randomly sampling reward weights $z$ and jointly learning a successor feature predictor $\psi_{z}$ and a policy $\pi_{z}$ conditioned on $z$. $\psi_{z}$ captures the successor feature of $\pi_{z}$, while $\pi_{z}$ is trained to maximize the reward described by $z$. FB [54, 55] follows the same paradigm but jointly learns a feature network by parameterizing the successor measure as an inner product between a forward and a backward representation. RaMP [9] removes the policy dependence of SF by predicting cumulative features from an initial state and an open-loop sequence of actions, which can be used for planning.
@@
6.3 Do DiSPOs enable zero-shot policy optimization across tasks?
@@
Figure 5: Transfer across tasks with DiSPOs and COMBO [63] in medium antmaze. Each tile corresponds to a different task, with color of the tile indicating the normalized return. DiSPOs successfully transfer across a majority of tasks, while MBRL [63] struggles on tasks that are further away from the initial location.
@@
We evaluate DiSPOs on transfer problems, where the dynamics are shared, but the reward functions vary. We train DiSPOs on the data distributions provided with the D4RL [15] and Hopper datasets. We identify the test-time reward by subsampling a small number of transitions from the offline dataset, relabeling them with the test-time rewards, and performing linear least squares regression. While DiSPOs in principle can identify the task reward from online experience, we evaluate in the offline setting to remove the confounding factor of exploration.
@@
Table 1 reports the episodic return on D4RL and Hopper tasks. DiSPOs are able to transfer to new tasks with no additional training, showing significantly higher performance than successor features (mismatch between training and evaluation policy sets), model-based RL (compounding error) and goal-conditioned RL (goal distribution misspecification). Notably, we show in Appendix E that DiSPOs are even competitive with goal-conditioned RL methods trained on test-time goals. The transferability of DiSPOs can also be seen in Fig  5, where we plot the performance of DiSPOs across various tasks (corresponding to different tiles in the maze). We see that DiSPOs have less degradation across tasks than model-based RL  [63].
@@
Although the DiSPO framework and theoretical results are derived under deterministic MDPs, we emphasize that the D4RL antmaze datasets are collected with action noise, emulating stochastic transitions. These results indicate that DiSPOs are practically applicable to some range of stochastic settings, although we expect it to perform better in purely deterministic settings.
6.4 Can DiSPOs solve tasks with arbitrary rewards?
@@
While methods like goal-conditioned RL [40, 17] are restricted to shortest path goal-reaching problems, DiSPOs are able to solve problems with arbitrary reward functions. This is crucial when the reward is not easily reduced to a particular “goal". To validate this, we evaluate DiSPOs on tasks that encode nontrivial human preferences in a reward function, such as particular path preferences in antmaze. In this case, we have different rewards that guide the agent specifically down the path to the left and the right, as shown in Fig 4. As we see in Table 3, DiSPOs and model-based RL obtain policies that respect human preferences and are performant for various rewards. Goal-conditioned algorithms are unable to disambiguate preferences and end up with some probability of taking each path.
@@
Table 2: Evaluation on non-goal-conditioned tasks. DiSPOs are able to solve non-goal-conditioned tasks, taking different paths in preference antmaze (Fig 4), while goal-conditioned RL cannot optimize for arbitrary rewards.
DiSPO (Ours) COMBO GC-IQL
@@
Table 3: Evaluation of trajectory stitching ability of DiSPOs. DiSPOs outperform non-stitching baselines, demonstrating their abilities to recombine outcomes across trajectory segments
DiSPO (Ours) RaMP DT
@@
6.5 Do DiSPOs perform trajectory stitching?
The ability to recover optimal behavior by combining suboptimal trajectories, or “trajectory stitching," is crucial to off-policy RL methods as it ensures data efficiency and avoids requirements for exponential data coverage. DiSPOs naturally enables this type of trajectory stitching via the distributional Bellman backup, recovering “best-in-data” policies for downstream tasks. To evaluate the ability of DiSPOs to perform trajectory stitching, we consider the environments introduced in  [46]. Here, the data only consists of trajectories that complete individual subtasks (e.g. grasping or placing), while the task of interest rewards the completion of both subtasks. Since the goal of this experiment is to evaluate stitching, not transfer, we choose the features as the task rewards $\phi(s)=r(s)$. We find that DiSPOs are able to show non-trivial success rates by stitching together subtrajectories. Since RaMP [9] predicts the summed features from a sequence of actions, and the optimal action sequence is not present in the dataset, it fails to solve any task. Likewise, return-conditioned supervised learning methods like Decision Transformer [10] do not stitch together trajectories and fails to learn meaningful behaviors.